OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−5, 5) and a directrix of y = -1.

OpenStudy (anonymous):

what do you think it is?

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

do they show answers or you have to type it in?

OpenStudy (anonymous):

f(x) = −one twelfth (x − 5)2 + 2 f(x) = one twelfth (x − 5)2 + 2 f(x) = −one twelfth (x + 5)2 + 2 f(x) = one twelfth (x + 5)2 + 2

OpenStudy (anonymous):

f(x) = −one twelfth (x − 5)2 + 2 f(x) = one twelfth (x − 5)2 + 2 f(x) = −one twelfth (x + 5)2 + 2 f(x) = one twelfth (x + 5)2 + 2

OpenStudy (anonymous):

@chefducky

OpenStudy (anonymous):

i dunno sorry

OpenStudy (anonymous):

3.09 lol.. I don't know the name of it...

OpenStudy (anonymous):

@mathmath333 @ParthKohli @paki @SolomonZelman @Secret-Ninja

OpenStudy (anonymous):

This should answer your question http://openstudy.com/study#/updates/52442865e4b0838bd44947f9

OpenStudy (anonymous):

it doesn't help... the answers there saying are wrong..